OpenStudy (anonymous):

Find all values of x which 5x^2-30x+2 less or equal to 7?

OpenStudy (anonymous):

I tried the derribative

OpenStudy (anonymous):

it did not work

OpenStudy (campbell_st):

so its \[5x^3 - 30x + 2\le 7\] subtract 7 from both sides... then divide by 5 then GQF

OpenStudy (anonymous):

This is an inequality. This is how we set this up:\[\bf 5x^2 -30x+2 \le 7 \implies 5x^2-30x-5 \le 0 \implies x^2 -6x - 1 \le 0\]Factor the equation and find the interval in which the parabola lies at and below 0.

OpenStudy (anonymous):

@luis0147896325

OpenStudy (anonymous):

Thank You @genius12

OpenStudy (anonymous):

wait i vant factor that

OpenStudy (anonymous):

@genius12 I cant